OpenStudy (anonymous):

'simple' algebra clarification within integration. hold, while i make the equations...

OpenStudy (anonymous):

\[\int\limits_{2}^{0} x^{2} \sqrt(1+x^{3} ) dx\] thats the original q and i have this picure: see attached.

OpenStudy (anonymous):

so my actual question is how does this guy get from the So... line to the next one, where u^{1/2}

OpenStudy (anonymous):

try substitution: u=1+x^3, du=3x^2

OpenStudy (anonymous):

i get the substitution but not the bit just after that

OpenStudy (anonymous):

that is, how does x squared suddenly equate to u to the power of one half?

OpenStudy (anonymous):

\[=\int\limits_{ }^{ }u ^{1/2}*1/3 du=2/9*u ^{3/2}=2/9 *(1+x^3)\]

OpenStudy (anonymous):

put the value of x -> 2to 0.... calculate

OpenStudy (anonymous):

are we good?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

whats wrong?

OpenStudy (anonymous):

after you let u= 1+x cubed then you've solved the derivative in terms of du how do you get the u to the power of one half?

OpenStudy (anonymous):

you have sq rt in original problem, don't you?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

i feel like your about to say something obvious

OpenStudy (anonymous):

ok. when you substitute \[u=1+x ^{3}, so: \sqrt{1+x ^{3}}=u ^{1/2}\]

OpenStudy (anonymous):

and x^2 will be take care off by du=x^2*1/3

OpenStudy (anonymous):

is it better now?

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

do it yourself on a peace of paper - always helps me :)

OpenStudy (anonymous):

i am doing it on paper :(

OpenStudy (anonymous):

i'll try to re-write differently...

OpenStudy (anonymous):

\[if u=1+x^{3}, then du=3x ^{2}dx\] \[means: x ^{2}dx=1/3 du\] put this substitution in original problem:

OpenStudy (anonymous):

=\[1/3\int\limits_{}^{}u ^{1/2}du=1/3 * 2/3 * u ^{3/2}=2/9 * (1+x ^{3})^{3/2}\] is it better?

OpenStudy (anonymous):

i got that far and subbed in sqrt (u) = sqrt(1+x cubed) so now i just have to \[\int\limits_{1}^{9} 1/3 u ^(1/2) du\]

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

im just not getting how your "one third, integral u to the one half wrt du = 1/3 * 2/3 *u^ 3/2

OpenStudy (anonymous):

i understand you can just "bring out" theh one third

OpenStudy (anonymous):

it's a number... I can put it in a front of integral any time

OpenStudy (anonymous):

yeah - thats fine.... but is the integral of u^1/2 = (2u/3)^3/2?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

integral = add 1 to the power & divide...

OpenStudy (anonymous):

oh ok so i typed it out wrong then, it should be = (3u/2)^3/2 * 1/3 from the front

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[=\int\limits_{}^{}u ^{1/2}=u ^{1/2+1}*1/(3/2)=...\]

OpenStudy (anonymous):

its cause a division of a fraction is the same as the multiplication of that fraction with the denominator and numerator reversed

OpenStudy (anonymous):

I meant: 1+1/2... right :)

OpenStudy (anonymous):

sigh

OpenStudy (anonymous):

i know... confusing sometime...

OpenStudy (anonymous):

remember the rules of integration: \[\int\limits_{}^{}x ^{n}dx=x ^{(n+1)}/(n+1)\]

OpenStudy (anonymous):

hm well does that mean that i should end in...

OpenStudy (anonymous):

18 as the answer?

OpenStudy (anonymous):

I got 56/9...let me check

OpenStudy (anonymous):

i went from [(2u^3/2)/3] from 2 to 0 = (2(1+x^3)^3/2)/3 =(2(1+2^3)^3/2)/3 =18

OpenStudy (anonymous):

oups... 52/9

OpenStudy (anonymous):

\[1/3 * 2/3 * (1+x ^{3})^{3/2}=2/9 (9)^{3/2}-2/9=...\] =6-2/9=52/9

OpenStudy (anonymous):

can you match/check the answer?

OpenStudy (anonymous):

how did you get 1/3 * 2/3 to be 2/9?

OpenStudy (anonymous):

(1*2)/(3*3)=2/9

OpenStudy (anonymous):

I use * as multiplication sign...confusing?

OpenStudy (anonymous):

using * as multiply isnt confusing.. just that a fraction divided by another fraction is the two numerators multiplied by each other, divided by the two denominators multiplied by each other..... how does that work?

OpenStudy (anonymous):

i just multiplied one third by 2 third... I'm lost... what is the question?

OpenStudy (anonymous):

i hate to tell you this, but its the crux of why i've failed at maths the last decade i dont understand division and multiplication of multiple fractions

OpenStudy (anonymous):

i've only just worked this out now.

OpenStudy (anonymous):

oups.... bummer!

OpenStudy (anonymous):

we had (1/3) : (3/2) =( 1/3)* (2/3)=2/9... I got to go, man :) Good Luck!!! you are on a right truck :))

OpenStudy (anonymous):

is : divide?

OpenStudy (anonymous):

no its times... ok thanks heaps

OpenStudy (anonymous):

: is division :) see ya!

OpenStudy (anonymous):

hm ok

OpenStudy (anonymous):

: is division * is multiplication sorry... gotta run

OpenStudy (anonymous):

ok... thanks again